• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name B. Craver

    Log entry time 20:27:44 on July 20, 2006

    Entry number 174609

    keyword=High Power Heater

    The target temperature reached 19.00 K and continued to slowly drop. It appears that the HPH is not turning on to regulate the temp - the PID output reads 400 W, but the readback reads a static 0.81. Setting the power manually has no effect on the readback value. The temperature has been stabilized by stepping the JT valve closed by a single step of 1.0. J.P. suggested rebooting the crate. This did not affect the readback. Next I reset the IOC using the web interface two separate times. Also, no effect. J.P suggests that the solution will probably require an entrance. Since, there are no plans to put beam on the target tonight, this entrance can be done tomorrow. I will simply maintain the target temperature stable with fine adjustments of the JT valve. Instructions on that in the next post.